OpenStudy (anonymous):

What is the fractional equivalent of 0.123123...?

OpenStudy (lollylau):

123/999 41/333

OpenStudy (lollylau):

Do you need proofs?

OpenStudy (anonymous):

mmh is it 41/333 ?

OpenStudy (lollylau):

a/1-r = sum of a + ar + ar^2, a is first term and r is ratio when r = 0.001 and a is 0.123, is becomes the first fraction.

OpenStudy (lollylau):

theres also another which i forgot.

OpenStudy (amistre64):

1000A =123.123123... - A = .123123... -------------------- 999A = 123 A = 123/999